OpenStudy (anonymous):

Translate to an equation. Thirteen less than six times a number is equal to 11 more than four times the number. What is the number?

OpenStudy (fibonaccichick666):

any ideas?

OpenStudy (anonymous):

no i doint know how to do this problem?

OpenStudy (anonymous):

This problem is asking you to translate English into mathematical symbols.

OpenStudy (anonymous):

ok but how do you do it?

OpenStudy (anonymous):

Do you know what mathematical symbols are?

OpenStudy (fibonaccichick666):

ok, so if I told you you had 5 dollars more than me, and I have 6 dollars, how many do you have

OpenStudy (fibonaccichick666):

could you write that out as an equation?

OpenStudy (anonymous):

"times" = x, etc...

OpenStudy (fibonaccichick666):

oh add on to that eq that the total between us is 17 dollars

OpenStudy (anonymous):

11

OpenStudy (anonymous):

i dont get it at all.

OpenStudy (anonymous):

Thirteen less than six times a number is equal to 11 more than four times the number. What is the number? We have the following mathematical elements in this sentence. "less than" is a difference, i.e., represented by a "-" "13" is a constant number. It is represented by itself. "13" "times" is multiplication, i.e., represented by an "x" or "*" "more than" is addition, i.e., represented by a "+" "four" is a number, i.e., represented by a "4" "a number" is a variable, i.e., represented by a letter, e.g., "a" Piece the equation together using those symbols.

OpenStudy (anonymous):

Typically when dealing with algebraic equations involving variables, we do not use "x" to represent multiplication, as it is the most often used variable in Algebra 1 and 2. Symbolically represent multiplication by using "*" or parenthesis.

OpenStudy (anonymous):

Print this out and study it: http://www.austincc.edu/tutor/students/resources/Translating%20English%20Words.pdf

OpenStudy (anonymous):

ok so 13-6*a=11+4*a right?

OpenStudy (anonymous):

Almost! It is 13 less than 6 times a number. So, 6*a - 13 = 4*a + 11

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

If you have 3 apples and I have 13 less than 6 * (the number of your apples), then I have 6 * (3 apples, which are your apples) - 13 apples total.

OpenStudy (anonymous):

That is your equation translated from English into mathematical symbols. You now need to solve for "a," which represents "a number."

OpenStudy (anonymous):

well can you show me how to do it then?

OpenStudy (anonymous):

Hint: get all your "a's" on one side of the equation.

OpenStudy (anonymous):

http://cstl.syr.edu/fipse/algebra/unit3/combine.htm Read up and show your attempt.